• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name Templon

    Log entry time 08:10:57 on June 5,2000

    Entry number 43158

    keyword=running shift summary

    Run 1792 ... 1.20 M, stopped due to problems with an RF cavity
    .
    beam down for about 1 hr with cavity problems leading ultimately to
    ... "resetting energy of linac"
    .
    Run 1793 ... 1.25 M
    ... during 1793, accelerator had a DAC fail and fast feedback was off for
    ... on the order of a few minutes. They say that they don't always catch
    ... it, so this should be a normal activity during shift (checking the feedback
    ... on the MEDM screen)
    .
    Run 1794 ... 1.25 M
    Run 1795 ... 1.25 M
    ...
    ... moving to pointing studies at 4:20, don't want to risk problems with
    ... beam at end of shift (hard cutoff of 8 AM).
    ..
    Run 1796 ......... 12C target, no raster, 5 uA.... 263 k events
    Run 1797 ..... 4 cm dummy, no raster, 5uA ... 230 k events
    Run 1798 ... 15 cm dummy, no raster, 5uA ... 288 k events
    Run 1799 ... 10 cm dummy, no raster, 5uA ... feedback was off, stop
    Run 1800 ... 10 cm dummy, no raster, 5uA ... 310 k events
    ... move back to carbon target to do HARP scan.
    .
    Run 1801 ... HARP scan. harp run number 926
    .. move back to LH2 target 15 cm to resume production run
    .
    Run 1802 ... production data, 42.5 uA ... 0.91 M events. Stopped by ROC14
    .................... disconnection.
    .
    Run 1803 ... 1.25 M
    Run 1804 ... 1.06 M
    .
    at 8 AM, MCC is not yet ready to take beam away, so we continue in
    principle. in practice they are having cavity problems, so at 8:10 AM
    no beam.